OpenStudy (anonymous):

If r=3+i and s=1-2i, express the following in the form a+bi, where a and b are real numbers. (g) r/s

OpenStudy (experimentx):

multiply by conjugate .. 1+2i on numerator and denominator

OpenStudy (anonymous):

r/s = (3+i)/(1-2i)=(3+i)(1+2i)/((1+2i)(1-2i))=3+7i-2 / 1+4 = (1+7i)/5 = 1/5 + 7i/5

OpenStudy (anonymous):

thanks :)